# LDAdam: Adaptive Optimization from Low-Dimensional Gradient Statistics

### v1

@@ Abstract
We introduce LDAdam [^1] Code available at: https://github.com/IST-DASLab/LDAdam, a memory-efficient optimizer for training large models, that performs adaptive optimization steps within lower dimensional subspaces, while consistently exploring the full parameter space during training. This strategy keeps the optimizer’s memory footprint to a fraction of the model size. LDAdam relies on a new projection-aware update rule for the optimizer states that allows for transitioning between subspaces, i.e., estimation of the statistics of the projected gradients. To mitigate the errors due to low-rank projection, LDAdam integrates a new generalized error feedback mechanism, which explicitly accounts for both gradient and optimizer state compression. We prove the convergence of LDAdam under standard assumptions, and show that LDAdam allows for accurate and efficient fine-tuning and pre-training of language models.
@@ 1 Introduction
When performing adaptive optimization within low-dimensional subspaces, a key challenge is to adapt the optimizer states to changes in the projection map. Adapting gradient statistics to new subspaces is not obvious and comes with two challenges. First, standard adaptive optimization algorithms are anisotropic: that is, the parameter-wise learning rate tuning implies that the choice of coordinate system used to express the gradient matrix impacts the optimization process. Second, the second-order statistics are nonlinear, which complicates estimation of the gradient statistics and error compensation. We address the first challenge by carefully selecting the coordinate system used for low-rank compression, i.e. the basis given by the most significant singular vectors. The second challenge is addressed by approximating the gradient statistics in the current subspace given the raw moments estimates in the previous subspace. Thus, in LDAdam, optimizer step is performed from the projection-aware intermediate optimizer states and the projected gradient, rather than from the optimizer states at the previous iteration and the full-rank gradient (Kingma & Ba, 2017). In turn, this leads to stronger convergence guarantees, and better practical performance.
The choice of projection map at each step is key to low-rank compression. Performing singular value decomposition (SVD) at each step would add prohibitive time overhead; instead, we make use of block power iteration with warm initialization (Bentbib & Kanber, 2015; Vogels et al., 2020), which can achieve a tight approximation of the most significant singular vectors of successive gradients, at low computational cost. To preserve the information from the last iterations kept within the optimizer states, we show that one can perform the block power iteration on a barycenter of the newly computed gradient and an average of the previous gradients. Finally, to mitigate the errors due to low-rank projection, we introduce a new variant of the error feedback mechanism (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) that supports compression of both gradient and optimizer states. Moreover, we propose a memory efficient implementation of it by reusing the space allocated to gradient accumulation.
@@ 2 Related work
Relative to GaLore, we propose the following new techniques: (1) we introduce a new projection-aware update rule (see Sections  3.1 and 3.2), which avoids accumulating low-rank gradients from different subspaces into the optimizer states, and performs frequent updates; (2) to correct for projection errors in both gradients and optimizer states, we introduce a new generalized error feedback mechanism (see Section 3.4). The convergence of GaLore is only guaranteed under a very strong “stable-rank” assumption, whereas we show convergence under standard conditions. Experimentally, we show that these improvements lead to consistently improved accuracy, and that they can be implemented efficiently in terms of both memory and runtime.
@@ 3.1 The Need for a Projection-Aware Update Rule
$\theta_{t+1}=\theta_{t}-\eta_{t}\bm{{\color[rgb]{0,0,1}\mathcal{U}_{t}}}\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+{\epsilon}}=\theta_{t}-\frac{\eta_{t}}{\sqrt{\bm{\hat{v}_{t}}}+{\epsilon}}\frac{1-\beta_{1}}{1-\beta_{1}^{t}}\sum_{\tau=1}^{t}\beta_{1}^{t-\tau}\bm{{\color[rgb]{0,0,1}\mathcal{U}_{t}}}\cdot\bm{\mathcal{U}_{\tau}^{\top}}\cdot g_{\tau}$ (1)
@@ 3.2 The LDAdam Update Rule
In LDAdam, we address two key challenges in low-rank updates: we introduce a projection-aware update rule that accounts for the change in the low-dimensional representation of the optimizer states, and we provide a generalized error feedback mechanism that accounts for the loss of gradient accuracy due to low-rank compression. Compared to standard adaptive optimization step, optimizer states are replaced by low-dimensional intermediate optimizer states, denoted by $m_{t-1/2}$, $v_{t-1/2}$; and the gradient is replaced by the accumulator $a_{t}=\mathcal{U}^{\top}_{t}\cdot A_{t}$ obtained by low-rank projection of the high-dimensional accumulator $A_{t}=g_{t}+\xi_{t}$, where $\xi_{t}$ is the (generalized) error feedback buffer, storing projection errors. In the following, we build the LDAdam algorithm step-by-step. The end-to-end construction is provided in Algorithm 1.
Our first objective is to find a mechanism to transfer gradient information from one lower dimensional vector space to another. For this, we take inspiration from orthogonal projections, which provide the best approximation in terms of $\ell_{2}$ error. The projection-aware update rule for the first moment estimate (Algorithm 1 Line 15) follows from the multiplication with the matrix $\mathcal{U}^{\top}_{t}\cdot\mathcal{U}_{t-1}$, which stands for both the projection matrix and the change of basis matrix. The matrix $\mathcal{U}^{\top}_{t}\cdot\mathcal{U}_{t-1}\in\mathbb{R}^{r\times r}$ is low-dimensional and thus allows efficient transition between subspaces, as no intermediate high-dimensional vector is created. We obtain the following update:
@@
Linearity of expectation leads to an autoregressive update rule for the first moment estimate (see Equation 4). With $(\langle\tilde{e_{i}},e_{j}\rangle)_{i,j}=(\mathcal{U}^{\top}_{t+1}\cdot\mathcal{U}_{t})_{i,j}$ the estimation of the first-order statistic in the coordinate system ~$\tilde{\mathcal{E}}_{t+1}$ resolves to an orthogonal projection when considering only the first $r$ components (i.e., after truncation to $i,j\leq r$), and the above discussion on compression of optimizer states remains valid. Following our statistical approach, we obtain the projection-aware update rule for the second moment estimates (Algorithm 1 Line 16) from the approximation below:
@@ 3.4 The Generalized Error Feedback Mechanism
14:         $a_{t}=\mathcal{U}^{\top}_{t}A_{t}$
15:         $m_{t-1/2}=\mathcal{U}^{\top}_{t}\mathcal{U}_{t-1}m_{t-1}$
16:         $v_{t-1/2}=(1-\beta_{2}^{t-1})\left|(\mathcal{U}^{\top}_{t}\mathcal{U}_{t-1})^{2}\cdot(\hat{v}_{t-1}-(\hat{m}_{t-1})^{2})+(\mathcal{U}^{\top}_{t}\mathcal{U}_{t-1}\cdot\hat{m}_{t-1})^{2}\right|$
17:    Optimizer states Adam-type update
@@ 4 Convergence Guarantees for LDAdam
All three assumptions are commonly used in adaptive optimization literature (Chen et al., 2019; Défossez et al., 2022; Li et al., 2022; Xie et al., 2023; Zhou et al., 2024). It is worth noting that the bounded gradient condition in Assumption 2 directly implies a bounded variance condition with a constant $\sigma^{2}=2G^{2}$. Therefore, for the purpose of asymptotic analysis, one can omit Assumption 3. However, we will show that in the non-convex convergence rate, the $\sigma^{2}$-term decays at a rate of $1/\sqrt{T}$, while all other $G$-terms decay at a faster rate of $1/T$. Theorem 1 states our general non-convex convergence result.
@@
5 LDAdam in Practice
5.1 Training Memory Footprint and Runtime
LDAdam’s memory savings over baseline Adam (Kingma & Ba, 2017) (see Table 1 for comparison) follow from the low dimensionality of the optimizer states: for gradients of shape $n\times m$ with $n\leq m$, the gradient statistics in a learning subspace of rank $r$ are stored as tensors of shape $r\times m$, and the additional projection matrix is of shape $n\times r$. As a result, the memory footprint of LDAdam’s optimizer states is only a fraction of the size of the model. For an $n\times n$ layer, the compression ratio of LDAdam relative to Adam is of $\frac{3}{2}\frac{r}{n}$. The Table 5 in the Appendix B reports peak memory for fine-tuning and pre-training tasks.
With the error buffer stored in the gradient variable, the error feedback mechanism is memory neutral. However, this comes at the cost of not supporting gradient clipping and per-layer weight updates (Lv et al., 2024b; Zhao et al., 2024). The latter allows memory savings proportional to the model size, but at the expense of gradient accumulation. Therefore, although applicable to most optimizers, it is not widely adopted, due to the large batch size used for large language model training (Touvron et al., 2023) and the memory-intensive nature of activations in attention layers.
Table 1: Optimizer comparison: parameter count during training for a weight layer of shape $n\times m$ with $n\leq m$ (i.e., left projection), training capabilities, and estimates of optimizer states memory footprint in half precision (models architecture is described in the Table 8).
Adam LDAdam GaLore
@@ 5.1 Training Memory Footprint and Runtime
Weights $nm$ $nm$ $nm$
Gradients $nm$ $nm$
Optimizer States $2nm$ $nr+2rm$ $nr+2rm$
@@
RoBERTa-base (r=8) 0.46 GB 0.15 GB 0.15 GB
Llama 350M (r=256) 1.37 GB 0.95 GB 0.95 GB
Llama-2 7B (r=32) 25.1 GB 1.22 GB 1.22 GB
Llama-2 7B (r=512) 25.1 GB 4.87 GB 4.87 GB
Runtime.
Compared to Adam, GaLore adds gradient compression; and compared to GaLore, LDAdam adds optimizer states projection-aware update and the generalized error feedback mechanism. Yet, the practical runtime differences are not substantial: concretely, on Llama-350M pre-training, LDAdam has a 10% longer runtime compared to GaLore, and is 15% slower than Adam. Table 6 in the Appendix B reports runtimes on fine-tuning and pre-training tasks.
5.2 Experiments
@@
Table 3: Results (accuracy) of fine-tuning Llama-2 7B on the GSM8K dataset.
@@
Across all fine-tuning experiments, LDAdam significantly outperforms GaLore, and is either on par or better than Adam. When ablating the LDAdam techniques, we observe that LDAdam without error feedback also outperforms GaLore, supporting LDAdam’s continuous subspace adaptation strategy over GaLore’s sequential update approach. The fact that we can match Adam once error feedback is enabled suggests that this technique is key to matching the practical performance of full-parameter training.
@@
Llama 130M 24.64 23.92 22.85 45.59 25.54 67.78 26.04
Llama 350M 18.08 18.35 17.31 - 19.86 - 20.03
Furthermore, we scale the experiments up to the Llama 1.3B parameter model, using the same hyperparameters but fixing the learning rate at $31e-$ 3 due to limited computational resources. Adam achieves a validation perplexity of 16.52, while LDAdam with rank r=16 achieves 15.81.
@@
Pre-training experiments also show that LDAdam has comparable performance to Adam, both in terms of convergence and validation accuracy, and confirm the improvements over GaLore. Further, for very high compression rates (i.e., r=8), LDAdam still converges, but its no-EF variant does not, nor does GaLore. This highlights the need for an error correction mechanism for pre-training tasks, which are known to require models updates of high rank. For lower compression (i.e. r=256), LDAdam outperforms the no-EF version and GaLore, which is correlated to their consistent projection errors. LDAdam even appears to slightly outperform Adam; we attribute this small improvement to the regularization effect of compression.
@@ 6 Discussion
In this work, we proposed a new low-rank learning method with the memory-efficient LDAdam optimizer, analyzed its convergence under standard assumptions, and provided empirical evidence of its ability to match the performance of Adam optimizer at a fraction of its memory cost.
Limitations. The LDAdam optimizer relies on Adam’s heuristic to estimate gradient statistics, but its development required the design of a specific intermediate optimizer state update rule. Therefore, efficient implementations of Adam are not directly applicable to LDAdam, and extending our work to other optimizers requires further work. On the experimental side, a natural next step would be to execute large-scale billion-parameter pre-training experiments. Doing so is feasible, but requires additional computational infrastructure and technical changes to the algorithm to allow for distributed training.
Impact Statement. This paper presents work that aims to advance the field of machine learning. We believe that memory-efficient optimization is a step toward democratizing large-scale model training, and thus provides opportunities to foster both the development of new applications and the research in the field. There are many societal concerns about the rapidly growing use of artificial intelligence, we feel that none of them is specifically related to our work.

### v4

@@ Abstract
We introduce LDAdam, a memory-efficient optimizer for training large models, that performs adaptive optimization steps within lower dimensional subspaces, while consistently exploring the full parameter space during training. This strategy keeps the optimizer’s memory footprint to a fraction of the model size. LDAdam relies on a new projection-aware update rule for the optimizer states that allows for transitioning between subspaces, i.e., estimation of the statistics of the projected gradients. To mitigate the errors due to low-rank projection, LDAdam integrates a new generalized error feedback mechanism, which explicitly accounts for both gradient and optimizer state compression. We prove the convergence of LDAdam under standard assumptions, and show that LDAdam allows for accurate and efficient fine-tuning and pre-training of language models. Code is available at https://github.com/IST-DASLab/LDAdam.
@@ 1 Introduction
When performing adaptive optimization within low-dimensional subspaces, a key challenge is to adapt the optimizer states to changes in the projection map. Adapting gradient statistics to new subspaces is not obvious and comes with two challenges. First, standard adaptive optimization algorithms are anisotropic: that is, the parameter-wise learning rate tuning implies that the choice of coordinate system used to express the gradient matrix impacts the optimization process. Second, the second-order statistics are nonlinear, which complicates estimation of the gradient statistics and error compensation. We address the first challenge by carefully selecting the coordinate system used for low-rank compression, i.e. the basis given by the most significant singular vectors. The second challenge is addressed by approximating the gradient statistics in the current subspace given the raw moments estimates in the previous subspace. Thus, with LDAdam, optimizer step is performed from the projection-aware intermediate optimizer states and the projected gradient, rather than from the optimizer states at the previous iteration and the full-rank gradient (Kingma & Ba, 2017). In turn, this leads to stronger convergence guarantees, and better practical performance.
The choice of projection map at each step is key to low-rank compression. Performing singular value decomposition (SVD) at each step would add prohibitive overhead; instead, we use block power iteration with warm initialization (Bentbib & Kanber, 2015; Vogels et al., 2020), which can achieve a tight approximation of the most significant singular vectors of successive gradients, at low cost. To preserve the information from the last iterations kept within the optimizer states, we show that one can perform the block power iteration on a barycenter of the newly computed gradient and an average of the previous gradients. Finally, to mitigate the errors due to low-rank projection, we introduce a new variant of the error feedback mechanism (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) that supports compression of both gradient and optimizer states. Moreover, we propose a memory efficient implementation of it by reusing the space allocated to gradient accumulation.
@@ 2 Related work
Relative to GaLore, we propose the following new techniques: (1) we introduce a new projection-aware update rule (see Sections 3.1 and 3.2), which avoids accumulating low-rank gradients from different subspaces into the optimizer states, and performs frequent updates; (2) to correct for projection errors in both gradients and optimizer states, we introduce a new generalized error feedback mechanism (see Section 3.4). The convergence of GaLore is only guaranteed under a very strong “stable-rank” assumption, whereas we show convergence under standard conditions. Experimentally, we show that these improvements lead to consistently improved accuracy, and that they can be implemented efficiently in terms of both memory and runtime.
@@ 3.1 The Need for a Projection-Aware Update Rule
$\theta_{t+1}=\theta_{t}-\eta_{t}\bm{{\color[rgb]{0,0,1}\mathcal{U}_{t}}}\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+{\epsilon}}=\theta_{t}-\frac{\eta_{t}}{\sqrt{\bm{\hat{v}_{t}}}+{\epsilon}}\frac{1-\beta_{1}}{1-\beta_{1}^{t}}\sum_{\tau=1}^{t}\beta_{1}^{t-\tau}\bm{{\color[rgb]{0,0,1}\mathcal{U}_{t}}}\cdot\bm{\mathcal{U}_{\tau}^{\top}}\cdot g_{\tau}.$ (1)
@@ 3.2 The LDAdam Update Rule
With LDAdam, we address two key challenges in low-rank updates: we introduce a projection-aware update rule that accounts for the change in the low-dimensional representation of the optimizer states, and we provide a generalized error feedback mechanism that accounts for the loss of gradient accuracy due to low-rank compression. Compared to standard adaptive optimization step, optimizer states are replaced by low-dimensional intermediate optimizer states, denoted by $m_{t-1/2}$, $v_{t-1/2}$; and the gradient is replaced by the accumulator $a_{t}=\mathcal{U}^{\top}_{t}\cdot A_{t}$ obtained by low-rank projection of the high-dimensional accumulator $A_{t}=g_{t}+\xi_{t}$, where $\xi_{t}$ is the (generalized) error feedback buffer, storing projection errors. In the following, we build the LDAdam algorithm step-by-step. The end-to-end construction is provided in Algorithm 1.
Our first objective is to find a mechanism to transfer gradient information from one lower dimensional vector space to another. For this, we take inspiration from orthogonal projections, which provide the best approximation in terms of $\ell_{2}$ error. The projection-aware update rule for the first moment estimate (Algorithm 1 Line 14) follows from the multiplication with the matrix $\mathcal{U}^{\top}_{t}\cdot\mathcal{U}_{t-1}$, which stands for both the projection matrix and the change of basis matrix. The matrix $\mathcal{U}^{\top}_{t}\cdot\mathcal{U}_{t-1}\in\mathbb{R}^{r\times r}$ is low-dimensional and thus allows efficient transition between subspaces, as no intermediate high-dimensional vector is created. We obtain the following update:
@@
Linearity of expectation leads to an autoregressive update rule for the first moment estimate (see Equation 4). With $(\langle\tilde{e_{i}},e_{j}\rangle)_{i,j}=(\mathcal{U}^{\top}_{t+1}\cdot\mathcal{U}_{t})_{i,j}$ the estimation of the first-order statistic in the coordinate system ~$\tilde{\mathcal{E}}_{t+1}$ resolves to an orthogonal projection when considering only the first $r$ components (i.e., after truncation to $i,j\leq r$), and the above discussion on compression of optimizer states remains valid. Following our statistical approach, we obtain the projection-aware update rule for the second moment estimates (Algorithm 1 Line 15) from the approximation below:
@@ 3.4 The Generalized Error Feedback Mechanism
14:         $m_{t-1/2}=\mathcal{U}^{\top}_{t}\mathcal{U}_{t-1}m_{t-1}$
15:         $v_{t-1/2}=(1-\beta_{2}^{t-1})\left|(\mathcal{U}^{\top}_{t}\mathcal{U}_{t-1})^{2}\cdot(\hat{v}_{t-1}-(\hat{m}_{t-1})^{2})+(\mathcal{U}^{\top}_{t}\mathcal{U}_{t-1}\cdot\hat{m}_{t-1})^{2}\right|$
16:    Optimizer states Adam-type update
17:         $a_{t}=\mathcal{U}^{\top}_{t}A_{t}$
@@ 4 Convergence Guarantees for LDAdam
All three assumptions, including the bounded gradient condition, are standard and commonly used in adaptive optimization literature (Reddi et al., 2019; Chen et al., 2019; Défossez et al., 2022; Modoranu et al., 2024). Attempting to relax the bounded gradient condition, some works (Shi et al., 2020; Zhang et al., 2022; Wang et al., 2024) resorted to using a strong growth condition (which is not necessarily a relaxation over bounded gradient condition) or bounded stochastic noise conditions (Li et al., 2023; Hong & Lin, 2024). Taniguchi et al. (2024) achieved an optimal rate by modifying Adam and only slightly relaxing the bounded gradient assumption by requiring the expected gradient norm to be bounded. The bounded gradient condition in Assumption 2 directly implies a bounded variance condition with a constant $\sigma^{2}=2G^{2}$. Therefore, for the purpose of asymptotic analysis, one can omit Assumption 3. However, we will show that in the non-convex convergence rate, the $\sigma^{2}$-term decays at a rate of $1/\sqrt{T}$, while all other $G$-terms decay at a faster rate of $1/T$. Theorem 1 states our general non-convex convergence result.
@@
5 Training Memory Footprint and Runtime
LDAdam’s memory savings over Adam (Kingma & Ba, 2017) (see Table 1 for comparison) follow from the low dimensionality of the optimizer states: for gradients of shape $n\times m$ with $n\leq m$, the gradient statistics in a learning subspace of rank $r$ are stored as tensors of shape $r\times m$, and the additional projection matrix is of shape $n\times r$. As a result, the memory footprint of LDAdam’s optimizer states is only a fraction of the size of the model. For an $n\times n$ layer, the compression ratio of LDAdam relative to Adam is of $\frac{3}{2}\frac{r}{n}$. Table 6 and Figure 3 reports peak memory for fine-tuning and pre-training tasks.
With the error buffer stored in the gradient variable, the error feedback mechanism is memory neutral. However, this comes at the cost of not supporting gradient clipping and per-layer weight updates (Lv et al., 2024b; Zhao et al., 2024). The latter allows memory savings proportional to the model size, but at the expense of gradient accumulation. To be precise, the standard GaLore algorithm does not support gradient accumulation (see Appendix D.2 for a detailed explanation), while its newer “retaining grad” version (Zhao et al., 2024) does, but at the expense of a memory overhead equal to the size of the model. Although applicable to most optimizers, per-layer weight updates is not widely adopted, due to the large batch size used for large language model training (Touvron et al., 2023) and the memory-intensive nature of activations in attention layers.
Table 1: Optimizer comparison: parameter count during training for a weight layer of shape $n\times m$ with $n\leq m$ (i.e., left projection), training capabilities, and estimates of optimizer states memory footprint in half precision (models architecture is described in Table 9).
Adam LDAdam GaLore (retaining grad) GaLore
@@ 5 Training Memory Footprint and Runtime
Weights $nm$ $nm$ $nm$ $nm$
Gradients $nm$ $nm$ $nm$
Optimizer States $2nm$ $nr+2rm$ $nr+2rm$ $nr+2rm$
@@
RoBERTa-base (r=8) 0.46 GB 0.15 GB 0.15 GB 0.15 GB
Llama 350M (r=256) 1.37 GB 0.95 GB 0.95 GB 0.95 GB
Llama-2 7B (r=32) 25.1 GB 1.22 GB 1.22 GB 1.22 GB
Llama-2 7B (r=512) 25.1 GB 4.87 GB 4.87 GB 4.87 GB
Compared to GaLore, LDAdam adds optimizer states projection-aware update and the generalized error feedback mechanism. Yet, the practical runtime differences are not substantial: on Llama-350M pre-training, LDAdam has a 10% longer runtime compared to GaLore, and is 15% slower than Adam. Table 7 and Figure 3 report runtimes, and Table 5 reports theoretical complexity in terms of standard matrix operations.
6 Experiments
@@
Table 3: Results of fine-tuning Llama-2 7B on the GSM8K dataset.
@@
Across all fine-tuning experiments, LDAdam outperforms GaLore, and is either on par or better than Adam. When ablating the LDAdam techniques, we observe that LDAdam without error feedback also outperforms GaLore, supporting LDAdam’s continuous subspace adaptation strategy over GaLore’s sequential update approach. The fact that we can match Adam once error feedback is enabled suggests that this technique is key.
@@
Llama 130M 24.64 23.82 22.65 45.17 25.32 67.78 26.04
Llama 350M 18.08 18.37 17.30 - 19.85 - 20.03
Furthermore, we scale the experiments up to the Llama 1.3B parameter model, using the same hyperparameters but fixing the learning rate at $45e-$ 4 due to limited computational resources. Adam achieves a validation perplexity of 14.86, while LDAdam with rank r=16 achieves 14.09.
@@
Pre-training experiments also show that LDAdam has comparable performance to Adam, both in terms of convergence and validation accuracy, and confirm the improvements over GaLore. Further, for very high compression rates (i.e., r=8), LDAdam still converges, but its no-EF variant does not, nor does GaLore. This highlights the need for an error correction mechanism for pre-training tasks, which are known to require models updates of high rank. For lower compression (i.e. r=256), LDAdam outperforms the no-EF version and GaLore, which is correlated to their consistent projection errors. For more details on the impact of the rank on training, see Figure 4. LDAdam even appears to slightly outperform Adam. We attribute this small improvement to compression, as sparse updates might enhance implicit regularization (Zhang et al., 2017; Neyshabur, 2017).
@@ 7 Discussion
We proposed a new low-rank learning method with the memory-efficient LDAdam optimizer, analyzed its convergence under standard assumptions, and provided empirical evidence of its ability to match the performance of Adam at a fraction of its memory cost. LDAdam relies on Adam to estimate gradient statistics, but its development required the design of a specific intermediate optimizer state update rule. Therefore, efficient implementations of Adam are not directly applicable to LDAdam, and extending our work to other optimizers requires further work. On the experimental side, a natural next step would be to execute large-scale billion-parameter pre-training experiments.
8 Ethics Statement
This paper presents work that aims to advance the field of machine learning. We believe that memory-efficient optimization is a step toward democratizing large-scale model training, and thus provides opportunities to foster both the development of new applications and the research in the field. There are many societal concerns about the rapidly growing use of artificial intelligence, we feel that none of them is specifically related to our work.